(y¹)² = (1-y²)/(1-x²); y² ½ when x =1


1
Expert's answer
2021-09-01T10:08:01-0400


\[ \begin{array}{l}
{\left(y'\right)}^{\mathrm{2}}\ \ =\ \ \frac{\left(\mathrm{1}-y^{\mathrm{2}}\right)}{\left(\mathrm{1}-x^{\mathrm{2}}\right)}\mathrm{;}\ \ \ \ \ \ y\left(\mathrm{1}\right)=\frac{\mathrm{1}}{\mathrm{2}}\ \ \ \ \ \ \ \ \left(accord\mathrm{in}g\ \ \ to\ \ the\ \ \ question\right) \\ 
 \\ 
y'\ =\ \ \sqrt{\frac{\left(\mathrm{1}-y^{\mathrm{2}}\right)}{\left(\mathrm{1}-x^{\mathrm{2}}\right)}}\  \\ 
 \\ 
\frac{dy}{\sqrt{\mathrm{1}-y^{\mathrm{2}}}}\ \ =\ \ \frac{dx}{\sqrt{\mathrm{1}-x^{\mathrm{2}}}}\ \ \ \left(By\ \ \ seperation\ \ of\ \ \mathrm{var}iables\right) \\ 
 \\ 
\int{\frac{dy}{\sqrt{\mathrm{1}-y^{\mathrm{2}}}}\ \ =\ \ \int{\frac{dx}{\sqrt{\mathrm{1}-x^{\mathrm{2}}}}}} \\ 
 \\ 
\mathrm{si}{\mathrm{n}}^{-\mathrm{1}}\left(y\right)\ \ =\ \ \ \mathrm{si}{\mathrm{n}}^{-\mathrm{1}}\left(x\right)\ \ +\ \ C \\ 
 \\ 
y=\mathrm{sin}\left(\mathrm{si}{\mathrm{n}}^{-\mathrm{1}}\left(x\right)\ \ +\ \ C\right) \\ 
 \\ 
from\ \ \ y\left(\mathrm{1}\right)=\frac{\mathrm{1}}{\mathrm{2}}\ \ ,\ \ \frac{\mathrm{1}}{\mathrm{2}}=\mathrm{sin}\left(\mathrm{si}{\mathrm{n}}^{-\mathrm{1}}\left(\mathrm{1}\right)\ \ +\ \ C\right) \\ 
 \\ 
\frac{\mathrm{1}}{\mathrm{2}}=\mathrm{sin}\left(\frac{\pi }{\mathrm{2}}\ +\ \ C\right)\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \left(\mathrm{Re}call\ \ that\ \ \ \mathrm{sin}\left(A+B\right)=\mathrm{sin}A\mathrm{cos}B{}{}\ \ +\ \ \mathrm{cos}A\mathrm{sin}B\right) \\ 
 \\ 
\frac{\mathrm{1}}{\mathrm{2}}=\mathrm{cos}\left(\ C\right),\ \ C=\mathrm{co}{\mathrm{s}}^{-\mathrm{1}}\left(\frac{\mathrm{1}}{\mathrm{2}}\right)\ =\ \ \frac{\pi }{\mathrm{3}} \\ 
 \\ 
y=\mathrm{sin}\left(\mathrm{si}{\mathrm{n}}^{-\mathrm{1}}\left(x\right)\ \ +\ \ \frac{\pi }{\mathrm{3}}\right) \end{array}
\] 




\end{document}